OpenStudy (anonymous):

I need Help What is the approximate volume of a cone with a height of 6 mm and radius of 18 mm? Use 3.14 to approximate pi, and express your final answer to the nearest tenth. mm3

OpenStudy (anonymous):

@camerondoherty @welshfella

OpenStudy (welshfella):

volume of a cone = (1/3)* 3.14*r^2*h where r = radius of the base and h = height

OpenStudy (welshfella):

so plug r = 18 mm and rh= 6 mm into this formula

OpenStudy (anonymous):

18*18 is 324

OpenStudy (welshfella):

* h not rh

OpenStudy (welshfella):

right now 324*6 = ?

OpenStudy (anonymous):

1944

OpenStudy (welshfella):

ok now multiply that by 1/3 then multiply the answer by 3.14

OpenStudy (anonymous):

2034.72

OpenStudy (welshfella):

right

OpenStudy (welshfella):

to the nearest 10th is 2014.7

OpenStudy (anonymous):

okay thanks

OpenStudy (welshfella):

* 2034.7

OpenStudy (anonymous):

ill give you a medal

OpenStudy (welshfella):

ty